# Navigating the Social Welfare Frontier: Portfolios for Multi-objective Reinforcement Learning

### v1

@@ Abstract
In many real-world applications of reinforcement learning (RL), deployed policies have varied impacts on different stakeholders, creating challenges in reaching consensus on how to effectively aggregate their preferences. Generalized $p$-means form a widely used class of social welfare functions for this purpose, with broad applications in fair resource allocation, AI alignment, and decision-making. This class includes well-known welfare functions such as Egalitarian, Nash, and Utilitarian welfare. However, selecting the appropriate social welfare function is challenging for decision-makers, as the structure and outcomes of optimal policies can be highly sensitive to the choice of $p$. To address this challenge, we study the concept of an $\alpha$-approximate portfolio in RL, a set of policies that are approximately optimal across the family of generalized $p$-means for all $p\in[-\infty,1]$. We propose algorithms to compute such portfolios and provide theoretical guarantees on the trade-offs among approximation factor, portfolio size, and computational efficiency. Experimental results on synthetic and real-world datasets demonstrate the effectiveness of our approach in summarizing the policy space induced by varying $p$ values, empowering decision-makers to navigate this landscape more effectively.
11footnotetext: Harvard University44footnotetext: Google Research55footnotetext: Equal contribution66footnotetext: Corresponding author, Email: cwkim@seas.harvard.edu
@@ 1 Introduction
This setting, which is often modeled using multi-objective reinforcement learning (MORL), arises in many RL applications, such as fair resource allocation in healthcare [32], cloud computing [27, 18] and communication networks [36, 7]. Recently, with the rise of large language models (LLMs), reinforcement learning from human feedback (RLHF) techniques that reflect the preferences of heterogeneous individuals have also been explored [6, 38, 26].
@@
Preference aggregation in such scenarios is often achieved by choosing a social welfare function, which takes the utilities of multiple stakeholders as input and outputs a scalar value representing the overall welfare [37, 9, 32, 13, 38, 26, 6]. However, selecting the appropriate social welfare function is a nontrivial task, as each function embodies a different notion of social welfare and can lead to vastly different outcomes for the involved stakeholders.
@@
In this work, we focus on a class of social welfare functions known as generalized $p$-means, a widely used class of social welfare functions in algorithmic fairness and social choice theory. Generalized $p$-means unify commonly used welfare functions such as Egalitarian welfare ($p=-\infty$), Nash welfare ($p=0$) and Utilitarian welfare ($p=1$), providing a smooth transition between these principles. Notably, this is known to be the only class of social welfare functions that satisfy several key axioms of social welfare, such as monotonicity, symmetry, and independence of scale [29, 22, 8, 24].
@@
However, each choice of $p$ represents a distinct notion of social welfare, creating a challenge in selecting the right $p$. In real-world applications, decision-makers must understand how policies change as $p$ varies and make informed decisions about which policy to deploy. Small changes in $p$ can sometimes lead to dramatically different policies, further complicating the selection process. Despite these challenges, much of the existing work assumes a fixed social welfare function — and hence a fixed value of $p$ — is given [19].
@@
To address this challenge, we propose a method for generating a finite portfolio of policies that approximately optimizes the $p$-means objectives across the entire spectrum of $p$ values. Our main algorithm, $p$-MeanPortfolio, sequentially selects finite $p$ values starting from $\infty$- ∞ to 1. These values are chosen so that the optimal policies at these points sufficiently cover the entire range of $p\in[-\infty,1]$ for a given approximation factor $\alpha$. We also propose a computationally efficient heuristic algorithm, which adaptively selects the next $p$ value from the intervals formed by previously chosen $p$ values.
@@
The portfolios provide a structured summary of approximately optimal policies, allowing decision-makers to navigate the implications of different $p$ values efficiently. For instance, the portfolio can illustrate how stakeholders’ outcomes evolve as $p$ varies, helping decision-makers make informed choices about which policy to deploy.
@@
Figure 1: Portfolio of three policies for the healthcare intervention setting (see Section 5 for details), obtained using Algorithm 1 with $\alpha=0.60$. The bar plots show total rewards induced by each policy across different education (left) and age (right) brackets. A “score” of $1.0$ corresponds to a baseline policy (not shown in the figure) used for comparison. The three policies impact various education and age brackets differently, offering a theoretically sound and diverse set of options for decision-makers.
@@ 1.1 Summary of Contributions
    We propose Algorithm $p$-MeanPortfolio (Algorithm 1) to compute a finite portfolio of policies that is $\alpha$-approximate for the generalized $p$-means objectives for any value of $p\in[-\infty,1]$. We provide theoretical guarantees, including an upper bound on the portfolio size and the number of MDPs solved to optimality, both expressed in terms of the approximation factor $\alpha$.
@@
    We evaluate our approach on three different domains, spanning synthetic to real-world problems. Our results show that a small portfolio can achieve near-optimal performance across all $p\in[-\infty,1]$. Moreover, the heuristic Budget-ConstrainedPortfolio constructs portfolios that closely match those produced by Algorithm $p$-MeanPortfolio, while significantly reducing computational costs.
@@ 1.2 Example
An illustrative example we consider in this work is a public health application, where the staff of a healthcare organization aims to train an intervention policy for multiple beneficiaries under a limited budget [32]. This setting can be captured by restless multi-armed bandits (RMABs) [35], which model sequential resource allocation problems under budget constraints. In this context, each beneficiary is associated with a state variable representing their level of engagement with the healthcare program, and the overall state of the problem is the concatenation of the states of all beneficiaries. At each time step, a policy determines which beneficiary to intervene on, subject to the budget constraint. Depending on the reward function used to train a policy, different socio-demographic groups among the beneficiaries can be prioritized. For example, one reward function might encode the preference to prioritize older populations, while another reward function might encode the preference to prioritize low-income populations. The decision-maker (e.g., healthcare organization staff) must train a policy that balances these conflicting objectives.
@@ 2 Related Work
Social Welfare Function and RL. In RL with multiple stakeholders, various social welfare functions have been explored, including generalized Gini-Welfare [37, 9], $p$-means [32, 13, 9], proportional-fairness [20], Nash welfare [21], among others. [2] investigated ordinal social welfare functions rather than cardinal ones.
@@
A related line of work has emerged in RLHF for LLMs, where social choice theory has been used to aggregate multiple reward models. For example, approaches leveraging Nash welfare [38], $\alpha$-fairness [26], and Egalitarian welfare [6] have been proposed. However, in both RLHF and broader multi-stakeholder RL, existing works typically assume a fixed social welfare function and focus on optimizing a single policy [19].
@@
Portfolios in MORL. In the broader MORL literature, the concept of portfolios of policies is well-established. A common approach is to compute solutions that approximate the Pareto front [25, 31, 28, 19]. While the Pareto front provides a characterization of trade-offs between conflicting objectives, it does not assume a specific set of social welfare functions. This generality, while powerful, presents challenges in practical multi-stakeholder settings. That is, the Pareto front does not directly correspond to any specific notion of welfare or utility, making it difficult to interpret the societal implications of these policies. Furthermore, without an explicit scalarization function to aggregate preferences, decision-makers must choose among Pareto-efficient policies without clear guidance on how to weigh the trade-offs, which is often critical in real-world applications where diverse preferences must be aggregated into a single deployable policy. Additionally, the Pareto front is typically large, making it prohibitively expensive to compute in practice [19].
A more refined notion of portfolio similar to the one we study is the concept of convex coverage sets. However, this concept only assumes weighted linear combinations of reward functions [30]. Our approach unifies diverse perspectives on $p$-means-based scalarization, bridging gaps left by existing methods that rely on a single fixed social welfare function or linear reward combinations.
@@
Portfolios in Optimization. In the optimization literature, [11] studied portfolios with stochastic guarantees for fixed objectives. Recent work in approximation algorithms also explores the notions of small-sized portfolios that approximately optimize a class of social welfare functions [16, 17, 14, 15, 5], by exploiting the combinatorial properties of facility location, scheduling and set cover problems. While these works operate in well-structured combinatorial settings, our work addresses the significantly more complex landscape of RL where the policy space is vast and computing optimal policies is inherently challenging. This necessitates novel algorithmic and theoretical advancements to efficiently construct portfolios with strong guarantees.
@@ 3.2 Social Welfare Function
We will use the following monotonicity property:
Lemma 3.2 (Theorem 1, Chapter 3.3.1 [4]).
For all strictly positive vectors $\mathbf{x}\in\mathbb{R}^{N}_{>0}$, and for all $p,q\in[-\infty,1]$ with $p<q$, we have $f(\mathbf{x},p)\leq f(\mathbf{x},q)$.
@@
In the MORL literature, $v^{(1)}(\cdot,p)$ is referred to as expected scalarized returns (ESR), while $v^{(2)}(\cdot,p)$ is known as scalarized expected returns (SER). In general, ESR is preferred when the expected total reward from a single execution is the primary focus, whereas SER is more suitable when policies are executed multiple times and rewards accumulate over iterations. Computing the optimal policy for $v^{(\ell)}(\cdot,p)$ requires specialized algorithms different from standard RL methods, which are not directly applicable. For example, [13] and [1] developed algorithms for $v^{(1)}(\cdot,p)$ and $v^{(2)}(\cdot,p)$, respectively. More detailed comparisons between these scalarization techniques and their solution algorithms are available in [1], [30], [28], [1], and [13]. While these aggregation rules are often studied independently in MORL, our work provides a unified framework applicable to both methods. We show that the proposed algorithms apply to both settings leveraging similar theoretical foundations.
@@ 3.3 Portfolio of Policies for All $p$-means
Given a $p\in[-\infty,1]$, aggregation rule $\ell\in\{1,2\}$, and an approximation factor $\alpha\in(0,1)$, a policy $\pi\in\Pi$ is an $\alpha$-approximation to the aggregation function $v^{(\ell)}(\cdot,p)$ if the value $v^{(\ell)}(\pi,p)$ achieved by $\pi$ is within factor $\alpha$ of the maximum achievable value of $v^{(\ell)}(\cdot,p)$ by policies in $\Pi$, that is,
@@
Given aggregation rule $\ell\in\{1,2\}$ and approximation factor $\alpha\in(0,1)$, a set of policies $\Pi^{\prime}\subseteq\Pi$ is called an $\alpha$-approximate portfolio for aggregation functions $v^{(\ell)}$ if for each $p\in[-\infty,1]$, there exists a policy in $\Pi^{\prime}$ that is an $\alpha$-approximation to $v^{(\ell)}(\cdot,p)$.
@@ 4.1 Algorithm $p$-MeanPortfolio
0:  $\alpha$-approximate portfolio $\Pi^{\prime}$ for the set of aggregation functions $v^{(\ell)}(\cdot,p),p\in[-\infty,1]$, assuming oracle access to solve problem (3)
@@
In this section, we introduce $p$-MeanPortfolio (Algorithm 1), which constructs an $\alpha$-approximate portfolio $\Pi^{\prime}$ for all $p\in[-\infty,1]$. In Theorem 4.1, we establish formal bounds on the portfolio size $|\Pi^{\prime}|$ and the number of oracle calls to solve (3). The full proof is provided in Appendix A.
@@
Given an MDP $\mathcal{M}$ with $N$ reward functions with condition number $\kappa$, set $\Pi$ of feasible policies, an aggregation rule $\ell\in\{1,2\}$, and a desired approximation factor $\alpha\in(0,1)$, Algorithm $p$-MeanPortfolio returns an $\alpha$-approximate portfolio of policies $\Pi^{\prime}$ for the set of aggregation functions $\{v^{(\ell)}(\cdot,p):p\in[-\infty,1]\}$. Further,
@@
Line search. The line search subroutine works as follows: given a $p\in[-\infty,1)$, it seeks to find some $b^{*}\geq p$ such that $\pi_{p}$ is an $\alpha$-approximation for all $q\in[p,b^{*}]$. To achieve this, it maintains lower and upper bounds $a,b$ on $b^{*}$ with $p\leq a<b\leq 1$ and iteratively refines these bounds.

### v2

@@ Abstract
In many real-world applications of reinforcement learning (RL), deployed policies have varied impacts on different stakeholders, creating challenges in reaching consensus on how to effectively aggregate their preferences. Generalized $p$-means form a widely used class of social welfare functions for this purpose, with broad applications in fair resource allocation, AI alignment, and decision-making. This class includes well-known welfare functions such as Egalitarian, Nash, and Utilitarian welfare. However, selecting the appropriate social welfare function is challenging for decision-makers, as the structure and outcomes of optimal policies can be highly sensitive to the choice of $p$. To address this challenge, we study the concept of an $\alpha$-approximate portfolio in RL, a set of policies that are approximately optimal across the family of generalized $p$-means for all $p\leq 1$. We propose algorithms to compute such portfolios and provide theoretical guarantees on the trade-offs among approximation factor, portfolio size, and computational efficiency. Experimental results on synthetic and real-world datasets demonstrate the effectiveness of our approach in summarizing the policy space induced by varying $p$ values, empowering decision-makers to navigate this landscape more effectively.
Machine Learning, ICML
@@ 1 Introduction
This setting, which is often modeled using multi-objective reinforcement learning (MORL), arises in many RL applications, such as fair resource allocation in healthcare (Verma et al., 2024a), cloud computing (Perez et al., 2009; Hao et al., 2023) and communication networks (Wu et al., 2018; Chen et al., 2021). Recently, with the rise of large language models (LLMs), reinforcement learning from human feedback (RLHF) techniques that reflect the preferences of heterogeneous individuals have also been explored (Chakraborty et al., 2024; Zhong et al., 2024; Park et al., 2024).
@@
Preference aggregation in such scenarios is often achieved by choosing a social welfare function, which takes the utilities of multiple stakeholders as input and outputs a scalar value representing the overall welfare (Yu et al., 2024; Cousins et al., 2024; Verma et al., 2024a; Fan et al., 2023; Zhong et al., 2024; Park et al., 2024; Chakraborty et al., 2024). However, selecting the appropriate social welfare function is a nontrivial task. Different functions encode distinct fairness criteria, and the resulting policies can lead to vastly different outcomes for stakeholders depending on the choice of the social welfare function.
@@
In this work, we focus on a class of social welfare functions known as generalized $p$-means, a widely used class of social welfare functions in algorithmic fairness and social choice theory. Each choice of $p$ represents a distinct notion of fairness, and the $p$-means unify commonly used welfare functions such as Egalitarian welfare ($p=-\infty$), Nash welfare ($p=0$) and Utilitarian welfare ($p=1$), providing a smooth transition between these principles. Notably, this is known to be the only class of social welfare functions that satisfy several key axioms of social welfare, such as monotonicity, symmetry, and independence of scale (Roberts, 1980; Moulin, 2003; Cousins, 2023; Pardeshi et al., 2024).
@@
In practice, the right choice of $p$ is often unclear in advance, and the decision-maker must understand how policies vary with $p$ in order to make informed choices about which $p$ (and thus which policy) to adopt. Small changes in $p$ can sometimes lead to dramatically different policies, and selecting a policy optimized for an arbitrary $p$ can lead to poor outcomes under a different $p$ value. Despite these challenges, much of the existing work assumes a fixed social welfare function — and hence a fixed value of $p$ — is given (Hayes et al., 2022).
@@
To address this challenge, we propose a method to compute a small yet representative set of policies that covers the entire spectrum of fairness criteria represented by $p\leq 1$. Our main algorithm, $p$-MeanPortfolio, sequentially selects finite $p$ values starting from $\infty$- ∞ to 1. These values are chosen so that the optimal policies at these points sufficiently cover the entire range of $p\leq 1$ for a given approximation factor $\alpha$. We also propose a computationally efficient heuristic algorithm, which adaptively selects the next $p$ value from the intervals formed by previously chosen $p$ values.
@@
The portfolios provide a structured summary of approximately optimal policies, allowing decision-makers to navigate the implications of different $p$ values efficiently. The decision-maker can review the impact of the choice of $p$ on the structure of policies, on relevant dimensions of interest, and make an informed choice about which policy to deploy.
@@
Figure 1: Portfolio of three policies for the healthcare intervention setting (see Section 5 for details), obtained using Algorithm 1 with $\alpha=0.60$. The bar plots show total rewards induced by each policy across different education (top) and age (bottom) brackets. A “score” of $1.0$ corresponds to a baseline policy (not shown in the figure) used for comparison. The three policies impact various education and age brackets differently, offering a theoretically sound and diverse set of options for decision-makers.
@@ 1.1 Summary of Contributions
    We propose Algorithm $p$-MeanPortfolio (Algorithm 1) to compute a finite portfolio of policies that is $\alpha$-approximate for the generalized $p$-means objectives for any value of $p\leq 1$. We provide theoretical guarantees, including an upper bound on the portfolio size and the number of MDPs solved to optimality, both expressed in terms of the approximation factor $\alpha$.
@@
    We evaluate our approach on three different domains, spanning synthetic to real-world problems. Our results show that a small portfolio can achieve near-optimal performance across all $p\leq 1$. Moreover, the heuristic Budget-ConstrainedPortfolio constructs portfolios that closely match those produced by Algorithm $p$-MeanPortfolio, while significantly reducing the computational cost.
@@ 1.2 Example
An illustrative example we consider in this work is a public health application, where the staff of a healthcare organization aims to train an intervention policy for multiple beneficiaries under a limited budget (Verma et al., 2024a). This setting can be captured by restless multi-armed bandits (RMABs) (Whittle, 1988), which model sequential resource allocation problems under budget constraints. In this context, each beneficiary is associated with a state variable representing their level of engagement with the healthcare program, and the overall state of the problem is the concatenation of the states of all beneficiaries. At each time step, a policy determines which beneficiary to intervene on, subject to the budget constraint. Depending on the reward function used to train a policy, different socio-demographic groups among the beneficiaries can be prioritized. For example, one reward function might encode the preference to prioritize older populations, while another reward function might encode the preference to prioritize low-income populations. The decision-maker (e.g., healthcare organization staff) must train a policy that balances these conflicting objectives.
@@ 2 Related Work
Social Welfare Function and RL. In RL with multiple stakeholders, various social welfare functions have been explored, including generalized Gini-Welfare (Yu et al., 2024; Cousins et al., 2024), $p$-means (Verma et al., 2024a; Fan et al., 2023; Cousins et al., 2024), proportional-fairness (Ju et al., 2024), Nash welfare (Mandal and Gan, 2023), among others. (Alamdari et al., 2024) investigated ordinal social welfare functions rather than cardinal ones.
@@
A related line of work has emerged in RLHF for LLMs, where social choice theory has been used to aggregate multiple reward models. For example, approaches leveraging Nash welfare (Zhong et al., 2024), $\alpha$-fairness (Park et al., 2024), and Egalitarian welfare (Chakraborty et al., 2024) have been proposed. However, in both RLHF and broader multi-stakeholder RL, existing works typically assume that a fixed social welfare function is given and focus on computing the corresponding optimal policy (Hayes et al., 2022). As a result, these methods do not offer guidance for settings where the appropriate choice of social welfare function is unclear in advance.
@@
Portfolios in MORL. In the broader MORL literature, the concept of portfolios of policies is well-established. A common approach is to compute solutions that approximate the Pareto front (Parisi et al., 2014; Van Moffaert and Nowé, 2014; Rădulescu et al., 2019; Hayes et al., 2022). While the Pareto front provides a characterization of trade-offs between conflicting objectives, it does not assume a specific set of social welfare functions. This generality, while powerful, presents challenges in practical multi-stakeholder settings. That is, the Pareto front does not directly correspond to any specific notion of welfare or utility, making it difficult to interpret the societal implications of these policies. Furthermore, without an explicit scalarization function to aggregate preferences, decision-makers must choose among Pareto-efficient policies without clear guidance on how to weigh the trade-offs, which is often critical in real-world applications where diverse preferences must be aggregated into a single deployable policy. Additionally, the Pareto front is typically large, making it prohibitively expensive to compute in practice (Hayes et al., 2022).
A more refined notion of portfolio similar to the one we study is the concept of convex coverage sets. However, this concept is limited to weighted linear combinations of reward functions (Roijers et al., 2013) and does not account for $p-$means. As a result, existing MORL methods, whether based on the Pareto frontier or weighted linear combinations, do not provide guarantees with respect to $p$-mean welfare functions. We include additional discussion comparing our method with several well-known MORL algorithms proposed in (Yang et al., 2019; Reymond et al., 2022; Alegre et al., 2023; Yu et al., 2024) in Appendix C.
@@
Portfolios in Optimization. In the optimization literature, (Drygala et al., 2024) studied portfolios with stochastic guarantees for fixed objectives. Recent work in approximation algorithms also explores the notions of small-sized portfolios that approximately optimize a class of social welfare functions (Gupta et al., 2023, 2025; Goel and Meyerson, 2006; Golovin et al., 2008; Chakrabarty and Swamy, 2019), by exploiting the combinatorial properties of facility location, scheduling and set cover problems. While these works operate in well-structured combinatorial settings, our work addresses significantly more complex landscape of RL where the policy space is vast and computing optimal policies is inherently challenging. This necessitates novel algorithmic and theoretical advancements to efficiently construct portfolios with strong guarantees.
@@ 3.2 Social Welfare Function
We also provide an illustrative example in Appendix B to provide the intuitive impact of various $p$-mean objectives. We will use the following monotonicity property:
Lemma 3.2 (Theorem 1, Chapter 3.3.1 (Bullen, 2003)).
For all strictly positive vectors $\mathbf{x}\in\mathbb{R}^{N}_{>0}$, and for all $p,q\leq 1$ with $p<q$, we have $f(\mathbf{x},p)\leq f(\mathbf{x},q)$.
@@
In the MORL literature, $v^{(1)}(\cdot,p)$ is referred to as expected scalarized returns (ESR), while $v^{(2)}(\cdot,p)$ is known as scalarized expected returns (SER). In general, ESR is preferred when the expected total reward from a single execution is the primary focus, whereas SER is more suitable when policies are executed multiple times and rewards accumulate over iterations. Computing the optimal policy for $v^{(\ell)}(\cdot,p)$ requires specialized algorithms different from standard RL methods, which are not directly applicable. For example, (Fan et al., 2023) and (Agarwal et al., 2022) developed algorithms for $v^{(1)}(\cdot,p)$ and $v^{(2)}(\cdot,p)$, respectively. More detailed comparisons between these scalarization techniques and their solution algorithms are available in Agarwal et al. (2022), Roijers et al. (2013), Rădulescu et al. (2019), Agarwal et al. (2022), and Fan et al. (2023). While these aggregation rules are often studied independently in MORL, our work provides a unified framework applicable to both methods. We show that the proposed algorithms apply to both settings leveraging similar theoretical foundations.
@@ 3.3 Portfolio of Policies for All $p$-means
Given a $p\leq 1$, aggregation rule $\ell\in\{1,2\}$, and an approximation factor $\alpha\in(0,1)$, a policy $\pi\in\Pi$ is an $\alpha$-approximation to the aggregation function $v^{(\ell)}(\cdot,p)$ if the value $v^{(\ell)}(\pi,p)$ achieved by $\pi$ is within factor $\alpha$ of the maximum achievable value of $v^{(\ell)}(\cdot,p)$ by policies in $\Pi$, that is,
@@
Given aggregation rule $\ell\in\{1,2\}$ and approximation factor $\alpha\in(0,1)$, a set of policies $\Pi^{\prime}\subseteq\Pi$ is called an $\alpha$-approximate portfolio for aggregation functions $v^{(\ell)}$ if for each $p\leq 1$, there exists a policy in $\Pi^{\prime}$ that is an $\alpha$-approximation to $v^{(\ell)}(\cdot,p)$.
@@ 4.1 Algorithm $p$-MeanPortfolio
0:  $\alpha$-approximate portfolio $\Pi^{\prime}$ for the set of aggregation functions $v^{(\ell)}(\cdot,p),p\leq 1$, assuming oracle access to solve problem (3)
@@
In this section, we introduce $p$-MeanPortfolio (Algorithm 1), which constructs an $\alpha$-approximate portfolio $\Pi^{\prime}$ for all $p\leq 1$. In Theorem 4.1, we establish formal bounds on the portfolio size $|\Pi^{\prime}|$ and the number of oracle calls to solve (3). The full proof is provided in Appendix A.
@@
Given an MDP $\mathcal{M}$ with $N$ reward functions with condition number $\kappa$, set $\Pi$ of feasible policies, an aggregation rule $\ell\in\{1,2\}$, and a desired approximation factor $\alpha\in(0,1)$, Algorithm $p$-MeanPortfolio returns an $\alpha$-approximate portfolio of policies $\Pi^{\prime}$ for the set of aggregation functions $\{v^{(\ell)}(\cdot,p):p\leq 1\}$. Further,
@@
Line search. The line search subroutine works as follows: given a $p\leq 1$, it seeks to find some $b^{*}\geq p$ such that $\pi_{p}$ is an $\alpha$-approximation for all $q\in[p,b^{*}]$. To achieve this, it maintains lower and upper bounds $a,b$ on $b^{*}$ with $p\leq a<b\leq 1$ and iteratively refines these bounds.